# Discrete-Time CRLB-based Power Allocation for CF MIMO-ISAC with Joint Localization and Velocity Sensing

### v2

@@ IV Power Allocation Algorithm for CF MIMO-ISAC
where the involved penalty sub-functions are given respectively by
@@
with ${f}_{l}({\bm{\rho}})\triangleq{\rm tr}\{{\bm{C}}_{\rm L}\}-\delta_{l}^{2}$ and ${f}_{v}({\bm{\rho}})\triangleq{\rm tr}\{{\bm{C}}_{\rm V}\}-\delta_{v}^{2}$. $q$ is the power of the penalty function. We consider $q=2$ to guarantee their continuous differentiability. Then, the objective function for the penalty problem can be written as
@@
where $\mu$ denotes the penalty factor.
Additionally, to enforce this equality constraint (22b) and reduce both computational complexity and the nonlinearity of the optimization problem, we adopt a projection operator instead of incorporating another penalty sub-function, i.e.,
@@
where ${\bm{\varpi}}\in\mathbb{R}^{N\times 1}$ is an arbitrary vector and ${\bm{\Theta}}\triangleq{\bm{I}}_{N\times N}-\frac{{\bm{1}}_{N}{\bm{1}}_{N}^{\rm T}}{N}$ is a projection matrix onto the zero space of ${\bm{1}}_{N}$. We have ${\bm{1}}_{N}^{\rm T}{\bm{\Theta}}={\bm{0}}_{N}^{\rm T}$. Thus, the equality constraint ${\bm{1}}_{N}^{\rm T}{\bm{\rho}}({\bm{\varpi}})=1$ in (22b) always holds by initializing ${\bm{\rho}}_{0}$ such that ${\bm{1}}_{N}^{\rm T}{\bm{\rho}}_{0}=1$. Consequently, ${\bm{\rho}}({\bm{\varpi}})$ is the sum of a constant PA vector ${\bm{\rho}}_{0}$ and a projection of ${\bm{\varpi}}$ onto the zero space of ${\bm{1}}_{N}$.
By substituting (25) into (24), we can convert the original constrained optimization to an unconstrained optimization problem as follows:
@@
Then, unconstrained optimization methods, e.g., CG and SD methods, can be employed to search for the optimal solution.
IV-A Modified Conjugate Gradient Algorithm
To numerically solve (26), we propose a projection-based MCG algorithm. The update at the $i$th iteration is given by
@@ IV-A Modified Conjugate Gradient Algorithm
where ${\bm{d}}_{i}$ is the search direction of $\mathcal{L}({\bm{\varpi}}_{i})$ and $\upsilon_{i}$ is the step size. The search direction using the MCG is updated by
@@
where $\varsigma_{i}$ denotes the deflection factor of a CG algorithm. Taking the derivative $\mathcal{L}({\bm{\varpi}})$ regarding ${\bm{\varpi}}$ yields
(ϖ)=bold-italic-$\small\nabla\mathcal{L}({\bm{\varpi}})=$ ${\bm{\Theta}}\nabla L({\bm{\rho}})={\bm{\Theta}}\big{[}{\bm{g}}+\nabla\alpha({\bm{\rho}})\big{]},$ (29)
@@
$\small\nabla\alpha({\bm{\rho}})=$ $\mu\big{(}\nabla\tilde{f}_{l}({\bm{\rho}})+\nabla\tilde{f}_{v}({\bm{\rho}})$
@@
Note that ${\bm{e}}_{n}\in\mathbb{R}^{N\times 1}$ is an index vector with its element at index $n$ being $1$ and the others being zeros.
Given the initial direction ${\bm{d}}_{1}=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{1})$, and using the idempotent property of the projection matrix, i.e., ${\bm{\Theta}}^{2}={\bm{\Theta}}$, we have ${\bm{\Theta}}{\bm{d}}_{i}={\bm{d}}_{i}$, $\forall{i>0}$. Thus, taking the projection of (27) leads to
@@
We have ${\bm{1}}_{N}^{\rm T}{\bm{d}}_{i}=0$ for $\forall{i>1}$ due to ${\bm{1}}_{N}^{\rm T}{\bm{d}}_{1}=0$ and ${\bm{1}}_{N}^{\rm T}{\bm{\Theta}}={\bm{0}}_{N}^{\rm T}$. Therefore, the equality constraint in (22b), that is, ${\bm{1}}_{N}^{\rm T}{\bm{\rho}}_{i}=1$, always holds at $i$th iteration for $\forall{i>1}$ if and only if the initial PA vector ${\bm{\rho}}_{1}$ satisfies ${\bm{1}}_{N}^{\rm T}{\bm{\rho}}_{1}=1$.
We now investigate the selection of the deflection factor $\varsigma_{i}$ and step size $\upsilon_{i}$ at each iteration, which is significant in guaranteeing the convergence of the MCG algorithm. One necessary condition for the convergence is the non-increasing objective function at each iteration, i.e.,
@@
To this end, the search direction ${\bm{d}}_{i}$ should be the descent direction and the step size $\upsilon_{i}$ should be correspondingly upper bounded by a threshold at each iteration.
IV-A1 Descent direction
To ensure a descent direction ${\bm{d}}_{i+1}$ at the $(i+1)$th iteration, it requires
@@ IV-A1 Descent direction
where the equality holds if and only if $\nabla L({\bm{\rho}}_{i+1})={\bm{0}}$. The initial direction ${\bm{d}}_{1}=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{1})$ is the descent direction of $L({\bm{\rho}}_{1})$ with $-\nabla L({\bm{\rho}}_{1})^{\rm T}{\bm{\Theta}}\nabla L({\bm{\rho}}_{1})\leq 0$ since the projection operator ${\bm{\Theta}}$ is a symmetric positive definite matrix. Substituting (33) into (35), we have
@@
where $\varsigma_{i}^{\rm b}$ is the bound for $\varsigma_{i}$ with its expression given by
@@
The numerator satisfies $\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})>0$ for any non-zero gradient $\nabla L({\bm{\rho}}_{i+1})$ due to the positive definite ${\bm{\Theta}}$.
When employing inexact line searches, the deflection selection proposed by Hestenes and Stiefel [40], referred to as HS method, for $\varsigma_{i}$ is preferable. Accordingly, we adopt the modified HS deflection, given by
$\small\varsigma_{i}^{\rm HS}=\dfrac{\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{q}}_{i}}{{\bm{d}}_{i}^{\rm T}{\bm{q}}_{i}},$ (38)
which is obtained from ${\bm{d}}_{i+1}^{\rm T}{\bm{q}}_{i}=0$ with ${\bm{q}}_{i}\triangleq{\bm{\Theta}}[\nabla L({\bm{\rho}}_{i+1})-\nabla L({\bm{\rho}}_{i})]$. In fact, when using the MCG algorithm, it is desirable to maintain orthogonality between the conjugate gradients of adjacent iterations, i.e.,
@@
(39) and (40) ensure the positivity of ${\bm{d}}_{i}^{\rm T}{\bm{q}}i$ and $\nabla L({\bm{\rho}}{i+1})^{\rm T}{\bm{q}}_{i}$, respectively. This allows for a positive conjugate gradient scaling factor $\varsigma_{i}^{\rm HS}$. Furthermore, (39) and (40) also imply that $|{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i+1})|\ll|{\bm{d}}_{i}^{\rm T}{\bm{q}}_{i}|$ and $|\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{q}}_{i}|\approx|\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})|$, respectively. As a result, $\varsigma_{i}^{\rm HS}$ is expected to be smaller than $|\varsigma_{i}^{\rm b}|$, thereby satisfying the deflection constraint defined in (36).
Rather than strictly relying on the potentially tight and unquantifiable conjugacy conditions in (39) and (40), we instead enforce the positivity of both the numerator and denominator in (38), along with the inequality $\varsigma_{i}^{\rm HS}<|\varsigma_{i}^{\rm b}|$. When these conditions hold, we set $\varsigma_{i}=\varsigma_{i}^{\rm HS}$. Otherwise, we reset the MCG iteration by letting $\varsigma_{i}=0$. To sum up, the deflection factor is updated by
${\varsigma_{i}=}\begin{cases}\varsigma_{i}^{\rm HS},\ \min\{\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{q}}_{i},{\bm{d}}_{i}^{\rm T}{\bm{q}}_{i},|\varsigma_{i}^{\rm b}|-\varsigma_{i}^{\rm HS}\}>0,\\ 0,\ {\rm else}.\end{cases}$ (41)
Note that, after every consecutive $N$ MCG steps with $\varsigma_{i}=\varsigma_{i}^{\rm HS}$, we reset $\varsigma_{i}=0$ to initiate a new cycle of MCG iterations. Such a reboot strategy can help mitigate the accumulation of numerical errors, restoring the orthogonality and conjugacy of direction vectors, and avoiding local minima in nonlinear and non-convex optimization.
IV-A2 Step size optimization
To optimize the step size selection, we define $f_{i}(\upsilon_{i})\triangleq L({\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i})$ with $f_{i}(0)=L({\bm{\rho}}_{i})$. The second-order Taylor approximation of $f_{i}(\upsilon_{i})$ at $\upsilon_{i}=0$ is given by
@@ IV-A2 Step size optimization
where $\nabla^{2}L({\bm{\rho}}_{i})$ is the Hessian matrix of $L({\bm{\rho}})$ at ${\bm{\rho}}_{i}$. According to (35), we have ${\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})\leq 0$. If $\|\nabla^{2}L({\bm{\rho}}_{i})\|$ is locally bounded, there exists a non-negative step size bound $\upsilon_{b}$ for given $\epsilon_{L}\in[0,1)$ such that
$\small f_{i}(\upsilon_{i})-f_{i}(0)\leq\epsilon_{L}\upsilon_{i}{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i}),\quad\text{for}\quad\upsilon_{i}\leq\upsilon_{b}.$ (43)
(43) always holds when ${\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}\leq 0$. In contrast, we have $\upsilon_{b}=\dfrac{2(\epsilon_{L}-1){\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})}{{\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}}$ when ${\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}>0$.
Due to the inherent high computational complexity of the Hessian matrix $\nabla^{2}L({\bm{\rho}}_{i})$ in calaulating $\upsilon_{b}$, we can apply an inexact line search (ILS) for the step size to enable
@@
According to (35), the term on the right-hand side of (44) is always non-positive. Thus, (44) can enable a descent function in (34). The ILS controlling factor $\epsilon_{L}$ is used to regulate the decrement of the function $L({\bm{\rho}})$ by moving ${\bm{\rho}}_{i}$ towards the direction ${\bm{d}}_{i}$ weighted by the step size $\upsilon_{i}$. When (44) is unmet for given $\upsilon_{i}$, the step size is scaled by
@@
where $\sigma_{\upsilon}\in(0,1)$ denotes the scaling factor and an initial step size is denoted as $\upsilon_{i}=\upsilon_{0,i}$.
Additionally, the penalty function does not inherently ensure that the individual power budget constraints at each transmitter are respected during the optimization process. As a result, negative PA values may arise in intermediate iterations. These infeasible PA values can lead to misleading CRLB evaluations, for instance, producing negative CRLBs instead of infinite ones, thereby falsely satisfying the CRLB constraints. To avoid unnecessary iterations within the infeasible region and to ensure physically meaningful results, it is crucial to enforce non-negative PAs at each iteration step. To this end, the step size $\upsilon_{i}$ needs to satisfy
@@
with $\rho_{n,i}$ denoting the $n$th element of ${\bm{\rho}}_{i}$ and similar for ${d}_{n,i}$. Thus, the initial step size for ILS can be selected as
$\small\upsilon_{0,i}=\epsilon_{\upsilon}\upsilon_{i}^{\rm b}$ (48)
with $\epsilon_{\upsilon}\in(0,1]$. Note that there must exist at least one component ${d}_{n,i}<0$ in each iteration $i$, unless ${\bm{d}}_{i}={\bm{0}}$, due to the constraint $\sum_{n=1}^{N}{d}_{n,i}=0$. The ILS method characterized in (44)-(48) can be seen as the extension and application of the Armijo’s Rule [41] in our work.
IV-A3 The algorithm implementation
A well-defined termination condition is essential for the proposed iterative algorithm. A natural choice, inspired by the ILS condition in (44), is to monitor the absolute change in the objective function between consecutive iterations, i.e., $|L({\bm{\rho}}_{i+1})-L({\bm{\rho}}_{i})|\leq\epsilon$, where $\epsilon$ is a predefined threshold with small value. To avoid numerical instability arising from the magnitude of the objective function itself, we apply a normalized termination condition, i.e., $|L({\bm{\rho}}_{i+1})-L({\bm{\rho}}_{i})|/|L({\bm{\rho}}_{i+1})|\leq\epsilon_{\rm th}$. This condition is employed in both the MCG-based and MSD-based algorithms.
We now present the complete PP-MCG-ILS. Generally, a large penalty factor may lead to numerical issues, such as an ill-conditioned Hessian matrix, which can degrade the convergence behavior of the MCG algorithm. To mitigate this, we adopt a sequential penalty strategy that gradually increases the penalty factor across successive iteration loops. Specifically, for a given penalty factor $\mu_{k}$, the MCG algorithm is executed until its termination condition is satisfied. Then, the penalty factor is updated as $\mu_{k+1}=\varphi\mu_{k}$ where $\varphi>1$ is a scaling constant, for the next loop of penalty-based iteration. This iterative refinement continues until the convergence criterion $\mu_{k}\alpha(\bm{\rho})\leq\epsilon_{\mu}$ is satisfied, where $\epsilon_{\mu}$ is a small, predefined threshold [39]. This sequential penalty strategy ensures both numerical stability and progressive enforcement of the constraint conditions. The detailed procedure of the proposed PP-MCG-ILS algorithm is summarized in Algorithm 1.

### v3

@@ IV-A Unconstrained Optimization Enabled by Penalty Function and Projection Operator
where the involved penalty sub-functions are
@@
with ${f}_{l}({\bm{\rho}})\triangleq{\rm tr}\{{\bm{C}}_{\rm L}\}-\delta_{l}^{2}$ and ${f}_{v}({\bm{\rho}})\triangleq{\rm tr}\{{\bm{C}}_{\rm V}\}-\delta_{v}^{2}$. $q$ denotes the power to the penalty function, which is set to $q=2$ in this work to ensure continuous differentiability. Then, the penalty-driven objective function is given by
@@
where $\mu$ denotes the penalty factor. Additionally, to reduce both computational complexity and the nonlinearity of the optimization problem, we adopt a projection operator to enforce the equality constraint (22b) instead of incorporating another penalty sub-function, i.e.,
@@
where ${\bm{\varpi}}\in\mathbb{R}^{N\times 1}$ is an arbitrary vector and ${\bm{\Theta}}\triangleq{\bm{I}}_{N\times N}-\frac{{\bm{1}}_{N}{\bm{1}}_{N}^{\rm T}}{N}$ is a orthogonal projection matrix onto the null space of ${\bm{1}}_{N}^{\rm T}$, i.e., ${\bm{1}}_{N}^{\rm T}{\bm{\Theta}}={\bm{0}}_{N}^{\rm T}$. Thus, the equality constraint ${\bm{1}}_{N}^{\rm T}{\bm{\rho}}({\bm{\varpi}})=1$ in (22b) always holds provided that ${\bm{1}}_{N}^{\rm T}{\bm{\rho}}_{0}=1$. Consequently, ${\bm{\rho}}({\bm{\varpi}})$ is the sum of a constant PA vector ${\bm{\rho}}_{0}$ and an orthogonal projection of $\bm{\varpi}$} onto the null space of ${\bm{1}}_{N}^{\rm T}$.
Substituting (25) into (24) can convert the original constrained optimization to an unconstrained problem as follows:
@@
Building upon the formulated problem, the next section presents unconstrained optimization methods for efficiently finding the PA solution.
IV-B Modified Conjugate Gradient Algorithm
To numerically solve (26), we employ an iterative search algorithm. The $(i+1)$th update is given by
@@ IV-B Modified Conjugate Gradient Algorithm
where ${\bm{d}}_{i}$ is the search direction of ${\bm{\varpi}}_{i}$ and $\upsilon_{i}$ is the step size. For second-order convergence, we consider using a CG algorithm, with the search direction updated by
@@
where $\varsigma_{i}$ denotes the deflection factor of a CG algorithm. $\nabla\mathcal{L}({\bm{\varpi}}_{i})$ is the derivative $\mathcal{L}({\bm{\varpi}})$ regarding $\bm{\varpi}$}, given by
$\small\nabla\mathcal{L}({\bm{\varpi}})=\$${\bm{\Theta}}\nabla L({\bm{\rho}})={\bm{\Theta}}\big[{\bm{g}}+\mu\nabla\alpha({\bm{\rho}})\big],$ (29)
@@
$\small\nabla\alpha({\bm{\rho}})=\$$\mu\big(\nabla\tilde{f}_{l}({\bm{\rho}})+\nabla\tilde{f}_{v}({\bm{\rho}})$
@@
Note that ${\bm{e}}_{n}\in\mathbb{R}^{N\times 1}$ is an index vector with its element at index $n$ being $1$ and the others being zeros. Different from classic CG algorithm, (29) is a projected gradient direction regarding $\bm{\rho}$}.
In the following, we will directly solve PA vector $\bm{\rho}$} instead of updating $\bm{\varpi}$}. By using the idempotent property of the projection matrix, i.e., ${\bm{\Theta}}^{2}={\bm{\Theta}}$, we firstly have ${\bm{\Theta}}\nabla\mathcal{L}({\bm{\varpi}})=\nabla\mathcal{L}({\bm{\varpi}})$ for the gradient $\nabla\mathcal{L}({\bm{\varpi}})$ in (29). Then, by setting the initial direction ${\bm{d}}_{0}=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{0})$ in (28), we further have ${\bm{\Theta}}{\bm{d}}_{i}={\bm{d}}_{i}$, $\forall{i>0}$. Finally, taking the projection of (27) leads to
@@
We refer to (32) and (33) as the penalty-function and projection-based modified CG method (PP-MCG). Different from (27), (32) yields the numerically iterative solution of the penalty-driven objective function (24). Additionally, with the search direction in (33), the equality constraint (22b), that is, ${\bm{1}}_{N}^{\rm T}{\bm{\rho}}_{i}=1$, always holds at the $i$th iteration for $\forall{i>0}$ if and only if the initial PA vector ${\bm{\rho}}_{0}$ satisfies ${\bm{1}}_{N}^{\rm T}{\bm{\rho}}_{0}=1$. This is because ${\bm{1}}_{N}^{\rm T}{\bm{d}}_{i}=0$ for $\forall{i>0}$ provided that ${\bm{1}}_{N}^{\rm T}{\bm{d}}_{0}=0$, due to ${\bm{1}}_{N}^{\rm T}{\bm{\Theta}}={\bm{0}}_{N}^{\rm T}$.
IV-C Parameter Optimization of PP-MCG Algorithm
A necessary condition for the convergence of the PP-MCG algorithm is that the objective function is non-increasing at each iteration, i.e.,
@@ IV-C Parameter Optimization of PP-MCG Algorithm
To this end, ${\bm{d}}_{i}$ should be a descent direction, and the step size $\upsilon_{i}$ should be optimized at each iteration. To ensure the descent direction, the deflection factor $\varsigma_{i}$ is crucial.
IV-C1 Descent direction
To ensure that ${\bm{d}}_{i+1}$ is a descent direction of $L({\bm{\rho}}_{i+1})$, it is required that
@@ IV-C1 Descent direction
where the equality holds if and only if $\nabla L({\bm{\rho}}_{i+1})={\bm{0}}$. The initial direction ${\bm{d}}_{0}=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{0})$ is descent with ${\bm{d}}_{0}^{\rm T}\nabla L({\bm{\rho}}_{0})\leq 0$ since the projection operator $\bm{\Theta}$} is a symmetric positive definite matrix. By substituting (33) into (35), the feasible range of the deflection factor can be obtained, i.e.,
@@
where $\varsigma_{i}^{\rm b}$ denotes the bound in selecting $\varsigma_{i}$, given by
@@
Note that $\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})>0$ for any nonzero gradient $\nabla L({\bm{\rho}}_{i+1})$, owing to the positive definiteness of $\bm{\Theta}$}.
When employing inexact line searches, the deflection selection for $\varsigma_{i}$ proposed by Hestenes and Stiefel [21], referred to as HS method, is preferable. Herein, we modify the HS deflection for the projection-driven iterative method, i.e.,
$\small\varsigma_{i}^{\rm HS}={\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{q}}_{i}}/{{\bm{d}}_{i}^{\rm T}{\bm{q}}_{i}},$ (38)
which is obtained from ${\bm{d}}_{i+1}^{\rm T}{\bm{q}}_{i}=0$ with ${\bm{q}}_{i}\triangleq{\bm{\Theta}}[\nabla L({\bm{\rho}}_{i+1})-\nabla L({\bm{\rho}}_{i})]$.
In fact, when using the MCG algorithm, it is desirable to maintain orthogonality between the conjugate gradients of adjacent iterations, i.e.,
@@
Note that (39) and (40) ensure the positivity of ${\bm{d}}_{i}^{\rm T}{\bm{q}}_{i}$ and $\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{q}}_{i}$, respectively. This allows for a positive conjugate gradient deflection factor $\varsigma_{i}^{\rm HS}$. Furthermore, (39) and (40) also imply that $|{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i+1})|\ll|{\bm{d}}_{i}^{\rm T}{\bm{q}}_{i}|$ and $|\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{q}}_{i}|\approx|\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})|$, respectively. As a result, $\varsigma_{i}^{\rm HS}$ is expected to be smaller than $|\varsigma_{i}^{\rm b}|$, thereby satisfying the deflection constraint defined in (36).
Rather than strictly relying on the potentially tight and unquantifiable conjugacy conditions in (39) and (40), we instead directly enforce the positivity of both the numerator and denominator in (38), along with the inequality $\varsigma_{i}^{\rm HS}<|\varsigma_{i}^{\rm b}|$. When these conditions hold, we set $\varsigma_{i}=\varsigma_{i}^{\rm HS}$. Otherwise, we reset the MCG iteration by letting $\varsigma_{i}=0$. To sum up, the deflection factor is updated by
$\small{\varsigma_{i}=}\begin{cases}\varsigma_{i}^{\rm HS},\ \min\{\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{q}}_{i},{\bm{d}}_{i}^{\rm T}{\bm{q}}_{i},|\varsigma_{i}^{\rm b}|-\varsigma_{i}^{\rm HS}\}>0,\\ 0,\ {\rm else}.\end{cases}$ (41)
Note that, after every consecutive $Q$ MCG steps with $\varsigma_{i}=\varsigma_{i}^{\rm HS}$, we reset $\varsigma_{i}=0$ to initiate a new cycle of MCG iterations. Such a reboot strategy can help mitigate the accumulation of numerical errors, restoring the orthogonality and conjugacy of direction vectors, and avoiding local minima in nonlinear and non-convex optimization.
IV-C2 Step size optimization
We have explored the selection of deflection factor above to ensure a descent direction in (35). The second step is to optimize the step size selection.
Let us define $f_{i}(\upsilon_{i})\triangleq L({\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i})$ with $f_{i}(0)=L({\bm{\rho}}_{i})$. The second-order Taylor approximation of $f_{i}(\upsilon_{i})$ at $\upsilon_{i}=0$ is given by
@@ IV-C2 Step size optimization
where $\nabla^{2}L({\bm{\rho}}_{i})$ is the Hessian matrix of $L({\bm{\rho}})$ at ${\bm{\rho}}_{i}$. Using the optimized deflection factor above, we have ${\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})\leq 0$. If $\|\nabla^{2}L({\bm{\rho}}_{i})\|$ is locally bounded, there will exist a non-negative step size bound $\upsilon_{i,\rm b}$ for a given constant $\epsilon_{L}\in[0,1)$ such that
$\small f_{i}(\upsilon_{i})-f_{i}(0)\leq\epsilon_{L}\upsilon_{i}{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i}),\quad\text{for}\quad\upsilon_{i}\leq\upsilon_{i,\rm b}.$ (43)
When ${\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}\leq 0$, (43) always holds. Otherwise, we have $\upsilon_{i,\rm b}=\dfrac{2(\epsilon_{L}-1){\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})}{{\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}}$ when ${\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}>0$. One can choose a step size at each iteration as $\upsilon_{i}=\delta_{i}\upsilon_{i,\rm b}$ for the latter case, with a small constant $\delta_{i}\in(0,1)$.
However, in this work, due to the high computational complexity of the Hessian matrix $\nabla^{2}L({\bm{\rho}}_{i})$ when computing $\upsilon_{i,\mathrm{b}}$, we adopt an inexact line search (ILS) strategy for step-size selection to satisfy (43), i.e.,
@@
Note that (44) can ensure a descent function in (34) since the term on the right-hand side is always non-positive. The ILS controlling factor $\epsilon_{L}$ is used to regulate the decrement of the function $L({\bm{\rho}})$ by moving ${\bm{\rho}}_{i}$ towards the direction ${\bm{d}}_{i}$ weighted by the step size $\upsilon_{i}$. When (44) is unmet for a given $\upsilon_{i}$, the step size is scaled by
@@
where $\sigma_{\upsilon}\in(0,1)$ is a scaling factor. An initial step size, denoted as $\upsilon_{i}=\upsilon_{i,0}$, can be chosen as described later.
Additionally, negative PA values may occur in intermediate iterations. Such infeasible values can lead to misleading CRLB evaluations, e.g., producing negative rather than infinite CRLBs, and may thus falsely appear to satisfy the CRLB constraints. To prevent unnecessary iterations within the infeasible region and to ensure physically meaningful results, it is essential to enforce non-negative PAs at each iteration step. To this end, the step size $\upsilon_{i}$ needs to satisfy
@@
with $\rho_{n,i}$ denoting the $n$th element of ${\bm{\rho}}_{i}$ and similar for ${d}_{n,i}$. Note that there must exist at least one component ${d}_{n,i}<0$ in each iteration $i$, unless ${\bm{d}}_{i}={\bm{0}}$, due to the constraint ${\bm{1}}_{N}^{\rm T}{\bm{d}}_{i}=0$. Thus, the initial step size for ILS can be selected as
$\small\upsilon_{i,0}=\epsilon_{\upsilon}\upsilon_{i}^{\rm b}$ (48)
with $\epsilon_{\upsilon}\in(0,1]$. The ILS method, characterized by (44)–(48), can be regarded as an adaptation of Armijo’s rule [7], customized to handle the non-negative PA optimization under projection constraint in our proposed framework.
IV-C3 The algorithm implementation
A well-defined termination condition is essential for the proposed iterative algorithm. A natural choice, inspired by the ILS condition in (44), is to monitor the absolute change in the objective function between consecutive iterations, i.e., $|L({\bm{\rho}}_{i+1})-L({\bm{\rho}}_{i})|\leq\epsilon$, where $\epsilon$ is a predefined small threshold. To avoid numerical instability arising from the magnitude of the objective function itself, we apply a normalized termination condition, i.e., $|L({\bm{\rho}}_{i+1})-L({\bm{\rho}}_{i})|/|L({\bm{\rho}}_{i+1})|\leq\epsilon_{\rm th}$. This is employed in both the MCG-based and the following MSD-based algorithms.
Finally, the selection of penalty factor is also important in guaranteeing a good learning performance. Generally, a large penalty factor may lead to numerical issues, such as an ill-conditioned Hessian matrix, which can degrade the convergence behavior of the MCG algorithm. To mitigate this, we adopt a sequential penalty strategy that gradually increases the penalty factor across successive iteration loops. Specifically, for a given penalty factor $\mu_{k}$, the MCG algorithm is executed until its termination condition is satisfied. Then, the penalty factor is updated as $\mu_{k+1}=\varphi\mu_{k}$ where $\varphi>1$ is a scaling constant, for the next loop of penalty-based iteration. This iterative refinement continues until the convergence criterion $\mu_{k}\alpha(\bm{\rho})\leq\epsilon_{\mu}$ is satisfied, where $\epsilon_{\mu}$ is a small, predefined threshold [32]. This sequential penalty strategy ensures both numerical stability and progressive enforcement of the constraint conditions. The detailed procedure of the PP-MCG-ILS algorithm is summarized in Algorithm 1.